OpenStudy (anonymous):

Subtract and simplify: (3x^2 + 7) – (x^2 – 6x + 4) I answered it, can someone check my answer? 2x^2+6x+3

hero (hero):

(3x^2 - x^2) + 6x + (7 - 4) (3 - 1)x^2 + 6x + 3 2x^2 + 6x + 3 Correct

hero (hero):

Looks like you're getting the hang of this :o)

OpenStudy (anonymous):

Hero?

OpenStudy (anonymous):

Wanna check another one of my answers?

hero (hero):

I can check your answers.

OpenStudy (anonymous):

k, here is another one

OpenStudy (anonymous):

Find the volume of the box: (x+3) (5x-4) (2x-1) I answered 10x^3+17x^2-35x+12

hero (hero):

(x + 3)(5x - 4)(2x - 1) = (x(5x - 4) + 3(5x - 4))(2x - 1) = (5x^2 - 4x + 15x - 12)(2x - 1) = (5x^2 + 15x - 4x - 12)(2x - 1) = (5x^2 + 11x - 12)(2x - 1) = 5x^2(2x - 1) + 11x(2x - 1) - 12(2x - 1) = 10x^3 - 5x^2 + 22x^2 - 11x - 24x + 12 = 10x^3 + (22 - 5)x^2 -(11 + 24)x + 12 = 10x^3 + 17x^2 - 35x + 12 Correct

OpenStudy (anonymous):

Great!

OpenStudy (anonymous):

You up for a few more?

hero (hero):

I guess

OpenStudy (anonymous):

Use FOIL to multiply the two binomials. (b + 3)(b – 9)

OpenStudy (anonymous):

i answered b^2-6b-27

hero (hero):

(b + 3)(b - 9) = b(b - 9) + 3(b - 9) = b^2 - 9b + 3b - 27 = b^2 -(9 - 3)b - 27 = b^2 - 6b - 27 Correct

OpenStudy (anonymous):

kk only like 2 more?

hero (hero):

In case you're wondering what that second step is, it's an application of the distributive rule. Highly recommend using it rather than FOIL

OpenStudy (anonymous):

10x(9x^2+60x+7) I answered 90x^3+60x^2+70x

OpenStudy (anonymous):

And ya i tried FOIL and I really didn't like it

hero (hero):

Actually if the original question is 10x(9x^2 + 60x + 7) then 90x^3 + 600x^2 + 70x is the correct result.

OpenStudy (anonymous):

Expand the squared binomial. (x – 4)^2 I didn't really understand this one but I guessed that it would be x^2-16?

OpenStudy (anonymous):

oops and the original was 6 not 60

hero (hero):

(x - 4)^2 = (x - 4)(x - 4) = x(x - 4) - 4(x - 4) = x^2 - 4x - 4x + 16 = x^2 -(4 + 4)x + 16 = x^2 - 8x + 16

OpenStudy (anonymous):

O ok

OpenStudy (anonymous):

The expression (2x + 3)(3x + 4) simplifies to which of the following?

OpenStudy (anonymous):

i don't understand that one ^^^

hero (hero):

Use distributive rule or FOIL: (2x + 3)(3x + 4) = 2x(3x + 4) + 3(3x + 4) = 6x^2 + 8x + 9x + 12 = 6x^2 +(8 + 9)x + 12 = 6x^2 + 17x + 12

hero (hero):

Simplify means the same as expand or FOIL in this case.

OpenStudy (anonymous):

Ok good because thats the one i chose!

OpenStudy (anonymous):

K thats all the questions thanks so much!!!!!

hero (hero):

Yup. I prefer distributive rule over FOIL because FOIL actually skips an intermediate step, which so happens to be the distributive rule.

hero (hero):

Furthermore, FOIL is just an acronym. Distributive rule is an actual mathematical property.

hero (hero):

Great job on all the questions.

hero (hero):

@skullpatrol, is there something I can help you with?

OpenStudy (skullpatrol):

@Hero No, I was just admiring your great work :-)

OpenStudy (anonymous):

@Hero just got all those questions graded I got 100%....Thnx!

hero (hero):

Awesome :D